OpenStudy (anonymous):

can anybody show me how to do this step by step? Simplify: √3(√12-√24)

OpenStudy (anonymous):

\(\sqrt{3*12}-\sqrt{3*24}=\sqrt{3^2*2^2}-\sqrt{3^2*2^2*2}=6-6\sqrt{2}\)

OpenStudy (anonymous):

where did the exponents come from?

OpenStudy (anonymous):

3*12=3*3*2*2

OpenStudy (anonymous):

oh so the factors of 12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the second 2?

OpenStudy (anonymous):

12=3*4=3*2*2

OpenStudy (anonymous):

24=3*8=3*2*2*2

OpenStudy (anonymous):

oh, ok.. see now, thank you :)

OpenStudy (anonymous):

yw